Acct 291

Salmone Company reported the following purchases and sales of its only product. Salmone uses a periodic inventory system. Determine the cost assigned to ending inventory using LIFO.
 

DateActivitiesUnits Acquired at CostUnits Sold at Retail
May 1Beginning Inventory310 units @ $16
5Purchase300 units @ $18
10Sales
220 units @ $26
15Purchase180 units @ $19
24Sales
170 units @ $27
Answer #1

In periodic inventory LIFO, the cost of goods sold and ending inventory is calculated at the end of the accounting period.

 

Total units available for sale = 310 + 300 + 180 = 790 units

 

Total units sold = 220 + 170 = 390 units

 

Ending inventory = Total units available for sale - Total units sold

= 790 – 390

= 400 units

 

Under the LIFO, latest units acquired become the cost of goods sold and oldest units will remain in inventory and reported as the cost assigned to ending inventory.


Date

Units

Rate

Total cost

May 1

310

$16

$4,960

May 5

90

$18

$1,620

Total

400


$6,580


∴ Cost assigned to ending inventory using LIFO = $6,580

answered by: Archie411
Answer #2
DateUnitsRateTotal cost
May 1          310 $           16 $               4,960
May 5            90 $           18 $               1,620
Total          400
 $               6,580


answered by: Archie411